Please explain in details

1

An investor has an initial margin requirement of 50% on his margin account and a maintenance margin requirement of 25%. The investor short sold 1,000 shares at $ 40 per share. What is the maximum price that the share can reach in the market before the investor receives a margin call?

2

The difference between mutual funds and hedge funds is There is no difference since there are both managed portfolios Hedge funds are not registered with the SEC since accepts only qualified investors such as high net worth individuals and institutional investors. None of the above.

3

The difference between mutual funds and Exchange traded funds (ETFs) is... ETFs shares are traded on exchanges all the time but mutual funds shares can be only traded at the end of the day when their net asset value is calculated. here is more variely of sharei ual funds None of the above
Answer #1

1. Total value of Shares = no of shares * Price = 1000* 40 = 40,000

Margin amount = 50%* 40,000 = 20,000

Margin Requirement = ( 1000* Share price - 20,000)/(1000* Share Price) = 25%
1000* Share Price - 20,000 = 250 * Share Price
Max Share Price to get margin call = 20,000/750 = 26.67

2. Option b is correct. Hedge funds are not required to be registered with the SEC whereas Mutual funds need to be registered.

3. Option a is correct. ETF are traded all the time but not on Net Asset value whereas Mutual Funds are traded on the absis of Net Asset Value.

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well